# Embodied Question Answering via Multi-LLM Systems

### v3

@@ 5.2.1 Baselines
Figure 5: LGX + CAM; Performance of CAM with $3$ independent agent that performed LGX for collecting observations. We use a 95%-5% train-test random split as in Figure 3. We experiment on Matterport environments with [LEFT] 215 nodes and [RIGHT] 53 nodes. Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy.
5.2.2 Results
We see the superior performance of CAM in Figure 3, where we run two sets of experiments, one to compare against MV and debating baseline in an $215$-node Matterport environment (left image) and the other in a $53$-node (right) environment. For each experiment, we report the mean and standard deviation accuracy calculated by Equation 1 over $5$ random seeds. We generate $600$ and $480$ queries for the $215$ and $53$-node environments, respectively, using the method described in Section 5.1. For both experiments, we train with $3$ independent answers to each query. The $3$ answers are from there LLM, where each time it answers it has only observations from one of the three rooms: bedroom, dining room, kitchen.
For the Central Model Training phase, the train-test random split of the generated queries and independent LLM-generated answers is $95$%-$5$%. In both environments, majority of the CAM method outperforms the MV and debating baselines. Specifically, XGBoost performs the best with $50$% $33$% higher accuracy than MV and debating, respectively. We also highlight the difference in time costs between the debate baseline and CAM. For each test question where three agents debate for $2$ rounds, the average inference time to make a final prediction is $30$ seconds, while the average time to train and run inference on the test set for XGBoost is $1.3$ seconds.
For a comparison on a larger test set with more input features using the $215$-node environment, in Figure 4 we have a random $80$%-$20$% train-test split. In this experiment, we have $7$ independent answers. Each answer is generated from the LLM, where the LLM only has access to the observations from a single room. The $7$ rooms are bedroom, kitchen, dining room, living room, family room, bathroom, and office. Each CAM method outperforms the baseline MV with XGBoost performing 46% higher.
We note that the two consistently lowest performing CAM methods and the ones closest to the MV and debating baselines are the linear models, Linear SVM and Logistic Regression Classifier. Intuitively, this similarity in performance makes sense as MV can be interpreted as a linear model with fixed equal weights, whereas Linear SVM and LR learn and adjust their weights based on the training data. Since we end up doing a majority vote on the final answers after the discussion, the debate baseline can also be construed as a linear model. These linear methods perform the worst suggesting the non-linearity of this EQA dataset.
6 CAM + Language-Guided Exploration
In this section, we evaluate our Multi-LLM framework by gathering observations using a SOTA exploration method. Rather than manually selecting which observations to provide to the LLM to answer the queries, we have all the LLM-based agents explore the household environment for a set amount of time before answering question.
@@ 6 CAM + Language-Guided Exploration
In the exploration phase, each subordinate agent is set in the Matterport house scan graph and performs an augmented variation of Language-Guided Exploration (LGX) as proposed by Dorbala et al. (2023). In traditional LGX, the LLM agent observes the different immediate household items around itself and chooses which one to navigate towards to find a specific target item, forming a chain of “hops” to the item.
@@
In our variation, we specifically provide the LLM with a system prompt that encourages the agent to explore as much of the house as possible. You are an embodied agent in a house. You want to aggressively explore the house and you want to find as many unique objects as you can.
@@
The agent then travels from node to node in the house graph. For each timestep, we document the timestep, node, and items observed. When the agent reaches a node we can then access the stored results and add them to the list of observed items. We have the $3$ agents perform $10$ steps of this augmented LGX and store the observations. The rest of the experiment process follows the same setup as described in Section 4 with the Question-Answering and Central Model Training phases. We perform both these experiments in the same two environments as in Section 5.2.2 and using a $95\%$-$5\%$ random train-test split over 5 seeds.
The performance of the various CAM methods and baselines in the larger $215$-node graph is shown on the left image of Figure 5 and the performance for the $53$-node environment is shown on the right. From both plots, we again see CAM model outperforms both baselines as with Section 5.2.2. In fact, with LGX for observation collection, we see that on average, all CAM methods have a higher test time accuracy than the non-learning aggregation baselines.
Interestingly, in the 53-node environment debating baselines do slightly worse on average than MV. This degradation occurs due to agents who were initially correct being swayed by incorrect agents during conversation. Overall, our Multi-LLM system can be utilized in conjunction with an LLM-based exploration method, emphasizing its practicality in real-world use cases. In the next section, we provide a feature importance of CAM.
7 Feature Importance Analysis
In the context of EQA, feature importance is a metric by which input impacts the model’s final answer. For our Multi-LLM system, feature importance can be used to examine if any agent response or query context (i.e. room or object) is relied heavily upon for training the central answer model. To examine which features hold the most weight in the final decision of CAM model, we look into studying the variable, or feature importance Breiman (2001); Fisher et al. (2018), of the CAM model. We measure feature importance using permutation feature importance (PFI). We define the PFI the $i$-th feature as,
@@ 7 Feature Importance Analysis
To examine feature importance, we train multiple DT-based CAM models, each with a different filtered query dataset from the $215$-node environment. In each training, the features comprise of the room and object name to represent the query and $4$ independent responses from LLMs that had observations either from the kitchen, dining room, bedroom, or office. The dataset however will only consist of rows with a given “Room” value. The model is trained and tested, and we then calculate feature importance using Equation 2.
@@
Table 3 summarizes our feature importance findings. The leftmost column specifies the room the dataset was filtered for, and the other columns show the average PFI of the object feature and independent responses for that filtered dataset over 5 trials. We see that in many cases, like when trained only on the dining room queries, the corresponding feature (i.e. response based only on dining room observation) does not have the highest PFI. We analyzed the offline LLM-generated answers based only on dining room observations and out of the $600$ responses, $549$ are “No”, or $91.5\%$. Thus, the CAM model has learned not to rely on these answers as much as the other responses even though they are based only on observations from the room in question. This outcome demonstrates how analyzing feature importance can help examine individual members of a multi-agent setup and further supports our approach utilizing independent agent responses.
8 Conclusion
@@
In this paper, we tackle Embodied Question Answering with a novel Multi-LLM framework that aggregates the independent responses of LLM-based agents and produces a final answer. To the best of our knowledge, we are the first to use multiple LLM-based agents for EQA. We propose a central answer model (CAM) that learns to weigh the responses of each agent to decide on a single answer to the given query. We show that various machine learning methods are suitable to train CAM and consistently outperform non-learning-based aggregation methods for ensemble outputs, specifically majority voting and multi-LLM debating. A key benefit of our system is that it requires no communication during inference, reducing time costs. Furthermore, our approach can be combined with SOTA methods for exploration in unknown households. We end this work by providing a feature importance analysis to study the effect of individual agents and target object specification on the final answer. We hope that this work will encourage further research in Multi-LLM approaches for EQA as LLM-based embodied agents become more household.
9 Limitations and Future Work
@@
This work has several limitations. First, the ground truth labels to train CAM can be difficult to obtain, especially if the household environment is constantly changing with non-stationary items. We leave it as future work to adapt this Multi-LLM setup for EQA in dynamic household environments. Second, our queries are all objective binary “Yes/No” questions. To better enhance practicality, future work should look into aggregation methods for situational, subjective, and non-binary questions. It would also be useful to remove the need to explicitly mention the target object in the query. Other areas directions for research could be then to apply our CAM aggregation method to question-answering tasks outside of Embodied AI such as long video understanding. Third, our results hinge on the quality of the observation data. In our work, we relied on the GLIP model. Future work should analyze the effect of scene understanding algorithms on our setup.
Appendix A Results on Additional Environments
@@
Below we provide additional results of our learning-based CAM methods against MV and debating on $7$ other Matterport environments. Similar to Figure 3, we use independent answers of three LLMs. One with observations only from dining room; another only with kitchen observations, and the last only with bedroom observations. Majority of the CAM methods consistently outperform the baseline majority vote (MV) and debate methods, with Decision Tree (DT) and XGBoost (XG) being the highest performing methods.
Figure 6: 20 nodes, 5 distinct rooms
Figure 7: 78 nodes, 11 distinct rooms
Figure 8: 111 nodes, 14 distinct rooms
Figure 9: 145 nodes, 15 distinct rooms
Figure 10: 114 nodes, 16 distinct rooms
Figure 11: 84 nodes, 13 distinct rooms
Figure 12: 43 nodes, 5 distinct rooms
Appendix B Decision Tree Visualization
@@
In this section, we showcase the potential interpretability of our learning-based CAM framework by visualizing the decision tree trained in Section 7. This tree in Figure 13 was trained on queries only pertaining to the kitchen and was trained on independent answers from four different LLM agents. One agent had only observations from kitchen; another only dining room; another bedroom; and the last one only had office observations. We see that the main feature used was actually the object query. We provide the object encoding labels in Table 4.
@@
Figure 13: Visualization of Decision Tree from Section 7.

### v4

@@ Introduction
Figure 3: LGX + CAM; Performance of CAM with $3$ independent agent that performed LGX for collecting observations. We use a 90%-10% train-test random split. We experiment on a Matterport environment with 215 nodes. We report the mean and standard deviation over $5$ trials. Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy.
Accuracy (%)
CAM-based Methods (Ours) Baselines
Environment Size (num of nodes) NN RF DT XG SVM SVM-L LR MV Debate
20 65.56 63.33 73.33 73.33 64.44 43.33 53.33 58.89 58.89
43 61.00 63.00 86.00 82.00 62.00 60.00 64.00 41.00 40.00
53 50.83 68.33 78.75 81.25 64.17 50.83 50.83 44.58 44.17
58 50.00 74.00 84.00 87.00 74.00 55.00 53.00 54.00 53.00
78 65.00 70.45 85.45 86.36 70.45 57.27 55.91 42.27 41.82
114 64.06 77.19 81.25 87.50 70.94 55.63 60.31 51.88 47.19
145 65.33 75.33 83.00 87.67 69.00 58.33 59.33 46.67 45.33
215 66.33 72.67 79.67 88.33 69.67 56.00 56.33 42.00 43.00
Accuracy (%) with Malicious Agent
CAM-based Methods (Ours) Baselines
Environment Size (num of nodes) NN RF DT XG SVM SVM-L LR MV Debate
20 61.11 62.22 73.33 73.33 64.44 44.44 53.33 61.11 58.89
53 60.42 68.33 78.75 81.25 64.17 49.58 50.83 46.67 44.17
145 68.00 75.33 83.00 87.67 69.00 59.33 59.33 47.67 45.67
215 64.33 73.67 79.67 88.33 69.67 56.33 56.33 43.67 43.00
Table 2: Performance measured in accuracy (Eq 1) of CAM with $3$ independent agent that performed LGX for collecting observations. We use a 90%-10% train-test random split. We experiment in $8$ different environments over $5$ trials. Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy. We repeated the experiment in $4$ of the environments but turned one of the agent malicious by providing responses opposite to what it believes are the correct answers. In cases where one agent is malicious, our CAM method is still better suited than MV and debate.
@@
We have all the LLM-based agents explore the household environment for a set amount of time before answering questions. In the exploration phase, each subordinate agent is set in the Matterport house scan graph and performs an augmented variation of Language-Guided Exploration (LGX) as proposed by Dorbala, Mullen Jr, and Manocha (2023). In traditional LGX, the LLM agent observes the immediate household items around itself and chooses which one to navigate to find a specific target item, forming a chain of “hops” to the item.
@@
In our variation, we instead provide the LLM with a system prompt that encourages the agent to explore as much of the house as possible. You are an embodied agent in a house. You want to aggressively explore the house and you want to find as many unique objects as you can.
The agent then travels from node to node in the house graph. For each timestep, we document the timestep, node, and items observed. When the agent reaches a node we can then access the stored results and add them to the list of observed items. We have the $3$ agents perform $10$ steps of this augmented LGX and store the observations. The rest of the experiment process follows the same setup as described in the Approach Section with the Question-Answering and Central Model Training phases. We use a $90\%$-$10\%$ random train-test split over $5$ seeds (0 to 4 are used as the seeds). Experiments are ran on an Apple M1 Pro and macOS 14.5.
Results: The performance of the various CAM methods and baselines in the $8$ graph environments is shown in Table 2. We test in environments of different sizes (measured in number of nodes) to showcase the generalization of the CAM approach across different size complexities. From the table, we see CAM model consistently outperforms both baselines in accuracy (Eq 1. We report the average accuracy over $5$ seeds. We see that the non-linear methods consistently achieve higher accuracy than the linear methods, suggesting that the dataset contains no hyperplane separating the queries with “Yes” and “No” ground-truth answers. Specifically, XGBoost attains the highest accuracy in all but one environment, suggesting that the boosted ensemble method can best fit the generated query-answer data.
Interestingly, in all but one of the environments, debating does slightly worse on average than MV. This degradation occurs due to agents who were initially correct being swayed by incorrect agents during conversation. Overall, our Multi-LLM system can be utilized in conjunction with an LLM-based exploration method, emphasizing its practicality in real-world use cases.
Method Agent 1 Agent 2 Agent 3
XG 38 $\pm$ 8.3 38 $\pm$ 7.0 38 $\pm$ 6.2
MV 97 $\pm$ 0.1 95 $\pm$ 0.1 96 $\pm$ 1.6
debate 97 $\pm$ 1.4 93 $\pm$ 3.1 91 $\pm$ 3.6
Table 3: Agreement (%) between each LGX agent and the final answers of 1) XGBoost-based CAM (XG), 2) Majority Vote (MV) and 3) debate in the $215$-node environment. We report the average and standard deviation over the $5$ trials. We see that XGBoost learns not to agree with any of the agents to achieve its high accuracy, suggesting more exploration or better object detection is needed.
Agreement: We analyze the agreement between each agent and the XGBoost model (the best performing CAM-based method), majority vote, and debate. We measure agreement between an agent and each central module method by the percentage of similar answers:
$Agr(s_{k},f)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{s_{k}^{i}=\hat{y}_{i}},$ (2)
where $i$ is index of the query parameterized by $(o,r)$. Table 3 reports the agreement over $5$ trials and we see that XGBoost has learned not to agree with the agents as much as MV and debate to attain its high accuracy. This low agreement could stem from the errors introduced by the object detection of GLIP and the question-answering of each individual agent. We further analyze the independent answers impact on training in the Feature Importance Analysis Section.
@@
Malicious Agent: We also show the benefits of CAM where one of the three agents is malicious. To turn the agent malicious we invert its independent answers so that all “Yes” answers turn to “No” and vice versa. We then train all CAM methods and execute baselines regularly. Thus, the malicious agent contributes an answer it believes to be incorrect. From Table 2, we see that our CAM methods still outperform the MV and debate baselines. We see that the debate baseline again does worse the majority vote, showing that the incorrect agent was able to sway the overall decision that was initially going to be correct. In the next section, we provide a feature importance of CAM.
Feature Importance Analysis
In the context of machine learning, feature importance is a metric by which input impacts the model’s final answer. For our Multi-LLM system, feature importance can be used to examine if any agent response or query context (i.e. room or object) is relied heavily upon for training the central answer model. To examine which features hold the most weight in the final decision of CAM model, we look into studying the variable, or feature importance (Breiman 2001; Fisher, Rudin, and Dominici 2018), of the CAM model. We measure feature importance using permutation feature importance (PFI). We define the PFI the $i$-th feature as,
@@
To examine feature importance, we train multiple DT-based CAM models, each with a different filtered query dataset from the $215$-node environment. To better highlight and focus on the feature importance aspect, we have an oracle exploration where we explicitly say which rooms and items an agent has in its observations memory when answering the questions. This feature importance analysis could be applied to dynamic agents that move around with LGX.
@@
In each training, the features comprise the room and object name to represent the query and $4$ independent responses from LLMs that had observations either from the kitchen, dining room, bedroom, or office. The dataset however will only consist of rows with a given “Room” value. The model is trained and tested, and we then calculate feature importance using Equation 3.
@@
Table 4 summarizes our feature importance findings. The leftmost column specifies the room the dataset was filtered for, and the other columns show the average PFI of the object feature and independent responses for that filtered dataset over 5 trials. We see that in many cases, like when trained only on the dining room queries, the corresponding feature (i.e. response based only on dining room observation) does not have the highest PFI. We analyzed the offline LLM-generated answers based only on dining room observations and out of the $600$ responses, $549$ are “No”, or $91.5\%$. Thus, the CAM model has learned not to rely on these answers as much as the other responses even though they are based only on observations from the room in question. This outcome demonstrates how analyzing feature importance can help examine individual members of a multi-agent setup for error analysis and further supports our approach utilizing independent responses. Please see the Appendix for a visualization of the Decision Tree to highlight the benefits of the interpretable ML-based CAM methods.
Conclusion and Limitations
@@
In this paper, we tackle a variation of Embodied Question Answering with a Multi-LLM framework that aggregates the independent responses of LLM-based agents and produces a final answer. While other methods have used a central module to aggregate independent agent answers for QA tasks, we specifically investigate this method in the context of embodied exploration in a household environment. We propose a central answer model (CAM) that learns to weigh the responses of each agent to decide on a single answer to the given query. We show that various machine learning methods are suitable to train CAM and consistently outperform non-learning-based aggregation methods for ensemble outputs, specifically majority voting and multi-LLM debating. A key benefit of our system is that it requires no communication during inference, reducing time costs. Furthermore, our approach can be combined with SOTA methods for exploration in unknown households. We also show that CAM is more robust to malicious agents than the majority vote and debate baslines. We end this work by providing a feature importance analysis to study the effect of individual agents and target object specification on the final answer. We hope that this work will encourage research in embodied Multi-LLM for QA as LLM-based agents become more household.
@@
This work has several limitations. First, the ground truth labels to train CAM can be difficult to obtain, especially if the household environment is constantly changing with non-stationary items. We leave it as future work to adapt this Multi-LLM setup for EQA in dynamic household environments. Second, our queries are all objective binary “Yes/No” questions. To better enhance practicality, future work should look into aggregation methods for situational, subjective, and non-binary questions.
Appendix A Appendix
Appendix B Code and Dataset
We provide an anonymous link to our code  [^1] https://anonymous.4open.science/r/CAM-embodied-llm-54CE/. The query dataset extracted is also within the submitted supplementary materials folder.
Appendix C Multi-LLM vs LLM
To motivate multiple LLM agents we provide a comparison between 1 agent and 3 agents in Table 5.
Environment Size Accuracy (%) of 1 Agent Accuracy (%) of 3 Agents (MV)
$53$ nodes $41.67$ $42.00$
$215$ node $44.17$ $44.58$
Table 5: Rise of accuracy with multiple dynamic agents
Appendix D Additional Visualization of Results
In this section, to focus on the aggregation abilities of CAM, we have an oracle exploration where we explicitly say which rooms and items an agent has in its observations memory when answering the questions.
@@ Appendix D Additional Visualization of Results
Below we provide additional visualizations of the results of our learning-based CAM methods against MV and debating on other Matterport environments. Like Table 2, we use independent answers of $3$ LLMs explored with LGX. Majority of the CAM methods consistently outperform the baseline majority vote (MV) and debate methods, with Decision Tree (DT) and XGBoost (XG) being the highest-performing methods.
Figure 4: 20 nodes
Figure 5: 43 nodes
Figure 6: 53 nodes
Figure 7: 58 nodes
Figure 8: 78 nodes
Figure 9: 114 nodes
Figure 10: 145 nodes
Appendix E Decision Tree Visualization
@@
In this section, we showcase the potential interpretability of our learning-based CAM framework by visualizing the decision tree trained in Section Feature Importance Analysis. This tree in Figure 11 was trained on queries only pertaining to the kitchen and was trained on independent answers from four different LLM agents. One agent had only observations from kitchen; another only dining room; another bedroom; and the last one only had office observations. We see that the main feature used was actually the object query. We provide the object encoding labels in Table 6.
@@
Figure 11: Visualization of Decision Tree from Section Feature Importance Analysis.